2(4z-1)=3(z+2) How do I solve this problem?
Answer by farohw(175)


2(4z-1) = 3(z+2)

First, multiply:
8z - 2 = 3z + 6

Gather like terms:
8z - 3z = 2 + 6

Then solve:
5z = 8

Divide both sides by 5 to isolate the variable, z:

5z/5 = 8/5

And, z = 8/5